# Improving Underwater Acoustic Classification Through Learnable Gabor Filter Convolution and Attention Mechanisms

### v1

@@ 5.1 Temporal proximity and dataset size evaluation
The results of the first experiment (a), which can be seen in Figure 6, support the idea that adjacent segments share sufficient similarities to facilitate classification. When the segments become more distant, the classification deteriorates, with the worst-case scenario showing a difference of 14 MCC percentage points. As the recording represents a moving target, performance variation is directly related to the change in physical distance between target and sensor, since the same target at different distances will not sound the same. Factors such as frequency attenuation and multi-path sound wave reflections can distort the perception of the received sound, thereby increasing the misclassification rates (stojanovic2008underwater). This result raises an important concern about the performance of Task 3. Since the classification performance is inaccurate when the recordings of the same vessels are separated in the temporal axis, the classification performance for different vessels will also be inaccurate, as is the case for Task 3. Additional improvements in signal processing could enhance the similarity of recordings produced under different environmental conditions and, consequently, improve the final classification of both Task 2 and Task 3.
@@
A series of ablation studies were conducted to evaluate the contribution of individual components within the GSE ResNeXt model, highlighting how classification performance is affected by the removal of each block. The results, shown in Table 4, demonstrate the effectiveness of incorporating the proposed components into the original architecture. Notably, performance was significantly impacted by the removal of the Gabor convolutional layer. The experiment using SE layers with traditional convolutions yielded the lowest performance. Since SE attention layers operate on channel-wise information, their effectiveness depends on the quality of the channel representations produced by the model. When traditional convolutions were used, the resulting channel representations were noisy, as illustrated in Figure 4, leading to reduced classification performance. When both the Gabor convolutional layers and SE attention mechanisms were removed, the classification exhibited less variation across folds. However, this configuration resulted in a substantial reduction in final classification coefficient than those that included Gabor convolutions.
Model MCC (%)
GSE ResNeXt 78.53 ± 2.50
- SE layer 78.03 ± 2.22
- Gabor Convolution 74.57 ± 2.73
- Gabor Conv, SE layer (ResNeXt) 76.83 ± 0.24
Table 4: Ablation studies performed with the GSE ResNeXt model for Task 2. First line represents the results using the complete model and subsequent lines represents the removal of each block.
@@ 6 Final Remarks
This research presents the GSE ResNeXt model, a residual neural network that incorporates Gabor convolutions in the initial layer and applies the squeeze-and-excitation attention mechanism in the network’s backbone. The combination of these two components functions as a two-dimensional frequency selection mechanism, enhancing both noise robustness and model convergence. The model’s performance in underwater acoustic classification was evaluated across three tasks of increasing difficulty, offering a comprehensive performance evaluation. The classification results demonstrate the superiority of the GSE ResNeXt model over commonly used architectures such as Xception and MobileNet V2, improving classification MCC by 2.58% and reducing training time by 28% when considering the scope of task 2.
@@
When assessing the impact of both Gabor convolutional layers and attention mechanisms on the results of Task 2, the contribution of these components is notable, showing a 2.2% improvement in classification performance. These results highlight the effectiveness of the frequency selection mechanism in classifying underwater acoustic targets, positioning it as a strong alternative for reducing training time and improving classification performance. Subsequent steps in this research could focus on mitigating the frequency attenuation inherent to the underwater acoustic domain, as well as evaluating the impact of the multi-path reflection of the sound wave on acoustic perception.

### v2

@@ 5.1 Temporal proximity and dataset size evaluation
The results of the first experiment (a), which can be seen in Figure 8, support the idea that adjacent segments share sufficient similarities to facilitate classification. When the segments become more distant, the classification deteriorates, with the worst-case scenario showing a difference of 14 MCC percentage points. As the recording represents a moving target, performance variation is directly related to the change in physical distance between target and sensor, since the same target at different distances will not sound the same. Factors such as frequency attenuation and multi-path sound wave reflections can distort the perception of the received sound, thereby increasing the misclassification rates [42]. This result raises an important concern about the performance of Task 3. Since the classification performance is inaccurate when the recordings of the same vessels are separated in the temporal axis, the classification performance for different vessels will also be inaccurate, as is the case for Task 3. Additional improvements in signal processing could enhance the similarity of recordings produced under different environmental conditions and, consequently, improve the final classification of both Task 2 and Task 3.
@@
5.2 Ablation Studies
A series of ablation studies were conducted to evaluate the contribution of individual components within the GSE ResNeXt model, highlighting how classification performance is affected by the removal of each block. To isolate the impact of the two core components that define the novelty of this approach, the Gabor module and the SE mechanism, the analysis focused on their incremental contribution. The remaining parts of the architecture follow standard ResNeXt design choices, which have been extensively validated in prior literature [51]. The results, shown in Table 5, demonstrate the effectiveness of incorporating these components into the original architecture for tasks 2 and 3.
Model MCC (%) Accuracy (%)
Task 2 Task 3 Task 2 Task 3
GSE ResNeXt 78.53 $\pm$ 2.50 64.86 $\pm$ 4.83 83.85 $\pm$ 1.85 73.31 $\pm$ 3.75
- SE layer 78.03 $\pm$ 2.22 61.86 $\pm$ 3.18 83.49 $\pm$ 1.66 71.17 $\pm$ 2.40
- Gabor Convolution 74.57 $\pm$ 2.73 65.46 $\pm$ 4.91 80.88 $\pm$ 2.03 73.89 $\pm$ 3.91
- Gabor Conv, SE layer (ResNeXt) 76.83 $\pm$ 0.24 53.00 $\pm$ 4.86 82.56 $\pm$ 0.22 64.09 $\pm$ 4.36
Table 5: Ablation studies performed with the GSE ResNeXt model for tasks 2 and 3. Data is provided as mean $\pm$ standard deviation across folds. First line represents the results using the complete model and subsequent lines represents the removal of each block.
In Task 2, performance was significantly impacted by the removal of the Gabor convolutional layer. The configuration using SE layers with traditional convolutions yielded the lowest performance. Since SE attention operates on channel-wise information, its effectiveness depends on the quality of channel representations produced by the model. With traditional convolutions, the resulting channel representations were noisy, as shown in Figure 6, leading to reduced classification performance. Removing both Gabor convolutions and SE attention produced less variation across folds, however, resulted in a substantial reduction in the final classification coefficient compared to configurations that included Gabor convolutions.
For Task 3, interpretation is limited by the considerable standard deviation between folds, which represents up to 9% of the average MCC. Still, the removal of both channel attention and Gabor convolutions affected the classification performance, with a reduction of 18% in MCC. Although removing Gabor convolutions achieved an average MCC about 0.6 percentage points higher than GSE ResNeXt, its performance varied greatly across folds, with a standard deviation of 4.91, making the difference not practically significant. Moreover, traditional convolutions showed slower and less stable convergence than Gabor convolutions, as illustrated in Figure 7, reinforcing their importance in the final model.
@@ 6 Final Remarks
This research presents the GSE ResNeXt model, a residual neural network that incorporates Gabor convolutions in the initial layer and applies the squeeze-and-excitation attention mechanism in the network’s backbone. The combination of these two components functions as a two-dimensional frequency selection mechanism, enhancing both noise robustness and model convergence. The model’s performance in underwater acoustic classification was evaluated across three tasks of increasing difficulty, offering a comprehensive performance evaluation. The classification results demonstrate the superiority of the GSE ResNeXt model over commonly used architectures such as Xception and MobileNet V2, improving classification MCC by 2.58% and reducing training time by 28% when considering the scope of Task 2 and 62% for Task 3.
@@
When assessing the impact of both Gabor convolutional layers and attention mechanisms on the results of Task 2, their contribution is evident, with a 2.2% improvement in classification performance, and even more pronounced for Task 3, which shows a 22.3% gain. These results highlight the effectiveness of the frequency selection mechanism in classifying underwater acoustic targets, positioning it as a strong alternative for reducing training time and improving classification performance.
Despite the improved classification performance, the proposed approach remains relatively heavy for embedded applications compared to lightweight models such as MobileNet V2 and MobileViT. This opens opportunities for compression or pruning to reduce computational cost. Efficiency could also benefit from reducing the input frequency range through resampling or band filtering, eliminating redundancies in spectral data. Subsequent steps in this research could focus on mitigating frequency attenuation inherent to the underwater acoustic domain, assessing the impact of multi-path sound wave reflections on acoustic perception, and performing cross-dataset validation by adapting the evaluation framework for smaller datasets.